# ET-Former: Efficient Triplane Deformable Attention for 3D Semantic Scene Completion From Monocular Camera

### v1

@@ III-C Triplane-based 3D Deformable Attention
Figure 6: Uncertainty Qualitative Results: In column 3, blue voxels show low uncertainty areas, while red voxels, highlighted by yellow rectangles, indicate high uncertainty. These uncertain areas are typically outside the camera’s FOV, or obscured by vegetation (row 1) or buildings (rows 2-3).
III-D Occupancy Prediction
In the first stage, the occupancy map $\mathbf{m}_{o}$ is estimated, which identifies occupied voxels in 3D space. As shown in Figure 4, we estimate depth from the RGB image, and then the depth pixels are projected into 3D space and voxelized into an occupancy map. The initial map is often inaccurate due to raw depth estimation. Following a similar approach to stage two, we use ResNet-50 and FPN to process image features, enhancing the occupancy map with the triplane-based deformable attention model described in Section III-C. In the triplane model, we use positional embeddings as the query features and aggregate them into three orthogonal planes. The output of the triplane model is an occupancy map $\mathbf{m}_{o}=\mathbf{l}_{o}([\mathbf{f}_{\text{cross}}^{l},\mathbf{f}_{\text{cross}}^{h},\mathbf{f}_{\text{cross}}^{d}])$, where $\mathbf{l}_{o}$ is a linear layer.
III-E Training Strategy
The two stages are trained separately, using a cosine scheduler. Given that the ground truth occupancy map $\mathbf{m}_{o}^{gt}$ is derived from a sequence of LIDAR data, which may include occlusions, we mask out missing and unlabeled voxels during training. Stage one uses a cross-entropy loss $\mathcal{L}_{ce}(\mathbf{m}_{o}^{gt},\mathbf{m}_{o})$ for occupancy prediction.
The second stage is trained in the formulation of CVAE, so we have the CVAE lower bound loss function:
@@ III-E Training Strategy
where $p()$ represents the probability. Besides $\mathcal{L}_{C}$, reconstruction losses including class-weighted cross-entropy loss $\mathcal{L}_{ce}$, and losses $\mathcal{L}_{scal}^{sem}$ and $\mathcal{L}_{scal}^{geo}$ from [cao2022monoscene] are used.
@@ IV Experiments
Dataset and Evaluation Metrics: In this work, various experiments were conducted to evaluate the proposed 3D uncertainty semantic prediction framework using the SemanticKITTI [behley2019semantickitti] dataset with monocular cameras. SemanticKITTI provides the semantic ground truth data in the range of $[0\ \text{m},51.2\ \text{m}]\times[-25.6\ \text{m},25.6\ \text{m}]\times[-2\ \text{m},4.4\ \text{m}]$. The semantic occupancy map has the shape of $256\times 256\times 32$ voxels, while the occupancy map for stage one has the shape $128\times 128\times 16$. Each voxel in the semantic occupancy map has 20 classes of semantic labels consisting of 19 semantic classes and 1 free class. Following the regular metric for SSC tasks, we utilize the Intersection over Union (IoU) to evaluate the occupancy prediction and the scene completion quality in stages one and two. For semantic evaluation, we use the mean IoU (mIoU) of the 19 semantic classes to evaluate the performance of the semantic prediction. Furthermore, we enumerate all 19 classes to evaluate the accuracy of the methods in each semantic class estimation. The training and evaluation are done by one GPU V100.
@@ IV-A Comparisons
Table I compares our approach, ET-Former, with state-of-the-art methods. Our model achieves the highest IoU and mIoU, surpassing other methods by at least $15.16\%$ in IoU and $18.12\%$ in mIoU. The triplane-based feature representation proves particularly effective for commonly seen, large semantic objects. We achieve the best semantic accuracy across multiple categories, such as road, building, car, vegetation, trunk, terrain, pole, and traffic-sign.
We evaluated the memory usage of ET-Former against VoxFormer [li2023voxformer] and others in training memory consumption (Table II). ET-Former reduces GPU memory usage by 3.7 GB $(25\%)$ compared to VoxFormer, and by 11.1 GB $(50.5\%)$ compared to Symphonize [jiang2024symphonize], while significantly outperforming them in IoU and mIoU. These results highlight the efficiency of the proposed method.
@@
Symphonize  [jiang2024symphonize] 41.44 13.44 22.0 Gb
@@
TABLE II: Memory Usage: Our method reduces GPU memory usage by 3.7 GB $(25\%)$ compared to VoxFormer [li2023voxformer], and by 11.1 GB $(50.5\%)$ compared to Symphonize [jiang2024symphonize], while outperforming them in IoU and mIoU.
For the evaluation of our first-stage occupancy prediction, we compared our results with those of VoxFormer, as presented in Table IV in Section VI. Our model achieves $14\%$ and $16.3\%$ improvements in IoU and Recall, respectively.
@@ IV-B Ablation Study
Table III presents an ablation analysis of the architectural components, using VoxFormer [li2023voxformer] as the baseline. Replacing VoxFormer’s second stage with our deterministic second stage model (where query features $\mathbf{f}_{o}$ are directly fed into the deformable triplane decoder instead of converting to Gaussian distributions as in (3)) improves IoU by 0.36 and mIoU by 0.28 compared to the baseline. Implementing our CVAE stochastic second stage further enhances performance by 0.63 IoU and 0.57 mIoU. Finally, combining our first stage model with the CVAE stochastic second stage significantly boosts performance by 7.57 IoU and 3.25 mIoU.
@@
TABLE III: Ablation Study: our key components (deformable triplane decoder, CVAE formulation, and proposed first stage model) each contribute to significant improvements in semantic estimation accuracy.
@@ IV-C Visualizations
Qualitative Results: Figure 5 qualitatively compares our approach with VoxFormer [li2023voxformer] and MonoScene [cao2022monoscene]. Blue rectangles highlight areas missed by other methods, while red rectangles indicate incorrect semantic estimations. VoxFormer shows considerable noise in its predictions, especially outside the camera’s FOV, and struggles with occluded areas (e.g., the driving road in the third row). MonoScene has difficulty estimating semantics in occluded areas and completing the map. In contrast, our ET-Former effectively handles occluded and out-of-FOV areas.
@@ V Conclusion, Limitations, and Future Work
ET-Former offers an efficient and robust solution for 3D semantic scene completion using a single monocular camera. It successfully achieves state of the art performance with much less memory usage, highlighting its potential for real-world applications in autonomous navigation systems. Our future work could focus on designing a more memory-efficient yet effective encoder. Additionally, we aim to design a navigation strategy that makes safe but also permissible decisions leveraging both semantic and uncertainty maps generated from the proposed algorithm.

### v2

@@ III-D Triplane-based 3D Deformable Attention
Figure 5: Uncertainty Qualitative Results: In column 4, blue voxels show low uncertainty areas, while red voxels, highlighted by yellow rectangles, indicate high uncertainty. These uncertain areas are typically outside the camera’s FOV, or obscured by vegetation (row 1) or buildings (rows 2-3).
III-E Occupancy Prediction
In stage 1, the occupancy map $\mathbf{m}_{o}$ is estimated, which identifies occupied voxels in 3D space. As shown in Figure 6, we estimate depth from the RGB image, and then the depth pixels are projected into 3D space, voxelized into a raw occupancy map, and finally converted to raw occupancy queries through positional embedding. The initial map is often inaccurate due to raw depth estimation. Following a similar approach to stage 2, we use ResNet-50 and FPN to process image features, enhancing the occupancy map with the triplane-based deformable attention model described in Section III-D. In the triplane model, we use raw occupancy query as the query features and aggregate them into three orthogonal planes. The output of the triplane model is an occupancy map $\mathbf{m}_{o}=\mathbf{l}_{o}([\mathbf{f}_{\text{cross}}^{l},\mathbf{f}_{\text{cross}}^{h},\mathbf{f}_{\text{cross}}^{d}])$, where $\mathbf{l}_{o}$ is a linear layer.
Figure 6: Stage 1: We employ the novel triplane-based deformable attention method in this stage. The estimated depth pixels are projected to the 3D space as raw occupancy queries to the triplane-based deformable model. The output of stage 1 is an occupancy map, the occupied voxels in which are then converted to occupancy queries.
III-F Training Strategy
The two stages are trained separately, using a cosine scheduler. Given that the ground truth occupancy map $\mathbf{m}_{o}^{gt}$ is derived from a sequence of LIDAR data, which may include occlusions, we mask out missing and unlabeled voxels during training. Stage 1 uses a cross-entropy loss $\mathcal{L}_{ce}(\mathbf{m}_{o}^{gt},\mathbf{m}_{o})$ for occupancy prediction.
Stage 2 is trained in the formulation of CVAE, so we have the CVAE lower bound loss function:
@@ III-F Training Strategy
where $p_{\theta}(\mathbf{z}|\mathbf{m}_{o},\mathbf{f}_{r})$ represents the approximate posterior distribution parameterized by the Voxel Encoder, and the KL divergence trains the distribution close to, $p(\mathbf{z})$, the Gaussian distribution of $\mathbf{z}$. Note that the posterior distribution $p_{\theta}(\mathbf{z}|\mathbf{m}_{o},\mathbf{f}_{r})$ was used to sample $\hat{\mathbf{f}}_{o}$ during training as Equation 3. $p(\mathbf{m}_{s}|\mathbf{f}_{r},\mathbf{z})$ represents the conditional distribution of generating the semantic map $\mathbf{m}_{s}$ from the latent sample $\mathbf{z}$ and image feature $\mathbf{f}_{r}$. Besides KL divergence, the reconstruction loss includes class-weighted cross-entropy loss $\mathcal{L}_{ce}$, and losses $\mathcal{L}_{scal}^{sem}$ and $\mathcal{L}_{scal}^{geo}$ from [cao2022monoscene].
@@ IV Experiments
Dataset and Evaluation Metrics: In this work, various experiments were conducted to evaluate the proposed 3D uncertainty semantic prediction framework using the SemanticKITTI [behley2019semantickitti] dataset with monocular cameras. SemanticKITTI provides the semantic ground truth data in the range of $[0\ \text{m},51.2\ \text{m}]\times[-25.6\ \text{m},25.6\ \text{m}]\times[-2\ \text{m},4.4\ \text{m}]$. The semantic occupancy map has the shape of $256\times 256\times 32$ voxels, while the occupancy map for stage 1 has the shape $128\times 128\times 16$. Each voxel in the semantic occupancy map has 20 classes of semantic labels consisting of 19 semantic classes and 1 free class. Following the regular metric for SSC tasks, we utilize the Intersection over Union (IoU) to evaluate the occupancy prediction and the scene completion quality in stages one and two. For semantic evaluation, we use the mean IoU (mIoU) of the 19 semantic classes to evaluate the performance of the semantic prediction. Furthermore, we enumerate all 19 classes to evaluate the accuracy of the methods in each semantic class estimation. The training and evaluation are done by one GPU V100.
@@ IV-A Comparisons
Table I compares our approach, ET-Former, with state-of-the-art methods. Our model achieves the highest IoU and mIoU, surpassing other methods by improving the SOTA scores of IoU by $15.16\%$ (from $44.71$ to $51.49$) and mIoU by $8.38\%$ (from $15.04$ to $16.30$). The triplane-based feature representation proves particularly effective for commonly seen, large semantic objects. We achieve the best semantic accuracy across multiple categories, such as building, car, other-vehicle, vegetation, trunk, terrain, and pole.
We evaluated the training memory usage of ET-Former against other approaches (Table II). ET-Former reduces GPU memory usage by 3.7 GB $(25\%)$ compared to VoxFormer, and by 11.1 GB $(50.5\%)$ compared to Symphonies [jiang2024symphonize], while significantly outperforming them in IoU and mIoU. These results highlight the efficiency of the proposed method.
@@
Symphonies  [jiang2024symphonize] 42.19 15.04 22.0 Gb
@@
TABLE II: Training Memory Usage On SemanticKITTI: Our method reduces GPU memory usage by 3.7 GB $(25\%)$ compared to VoxFormer [li2023voxformer], and by 11.1 GB $(50.5\%)$ compared to Symphonies [jiang2024symphonize], while outperforming them in IoU and mIoU.
To evaluate our stage 1 occupancy prediction model proposed in Section III-E, we conducted a comparative analysis with VoxFormer [li2023voxformer]. To ensure a fair comparison, we utilize the same depth estimation method, MobileStereoNet [shamsafar2022mobilestereonet], as employed by the stage 1 of VoxFormer. Subsequently, as depicted in Figure 6, we apply our triplane-based method to generate the occupancy map and queries. The quality of the occupancy map predictions is compared between our stage 1 approach and VoxFormer’s, as shown in Table III. Our approach shows substantial improvements, with a $14.2\%$ increase in IoU and a $16.3\%$ increase in Recall, resulting in a more accurate occupancy map and more efficient occupancy queries. Figure 7 shows the qualitative comparison between our stage 1 approach and VoxFormer’s. We observe that our approach, ET-Former, can estimate details more accurately and generate more complete occupancy map than VoxFormer. The improvement in stage 1 ultimately contributes to the performance improvement of the subsequent semantic predictions, as shown in Table IV.
Methods IoU Recall
VoxFormer’s stage 1 52.03 61.52
Our stage 1 59.41 71.56
TABLE III: Results of Stage 1 On SemanticKITTI test [behley2019semantickitti]: Compared with the VoxFormer’s stage 1 model, our stage 1 model achieves $14.2\%$ and $16.3\%$ improvements in IoU and Recall values, respectively.
RGB
Refer to caption Refer to caption
VoxFormer [li2023voxformer]
Refer to caption Refer to caption
ET-Former
Refer to caption Refer to caption
Figure 7: Stage 1 Restuls: The major differences between the stage 1 results of ET-Former and VoxFormer are marked by red rectangles. In the first column, ET-Former demonstrates superior performance in detecting trees and cars, providing more detailed and accurate representations. The second column illustrates ET-Former’s ability to generate a more comprehensive map, including the detection of distant roads that VoxFormer has missed.
@@ IV-B Ablation Study
Table IV presents an ablation analysis of the architectural components, using VoxFormer [li2023voxformer] as the baseline. Replacing VoxFormer’s stage 2 with our deterministic stage 2 model (where query features $\mathbf{f}_{o}$ are directly fed into the deformable triplane decoder instead of converting to Gaussian samples $\hat{\mathbf{f}}_{o}$ as in Equation 3) improves IoU by 0.36 and mIoU by 0.28 compared to the baseline. Implementing our CVAE stochastic stage 2 further improves performance by 0.63 IoU and 0.57 mIoU. Finally, combining our stage 1 model with the CVAE stochastic stage 2 significantly boosts performance by 7.57 IoU and 3.25 mIoU.
@@
TABLE IV: Ablation Study On SemanticKITTI test [behley2019semantickitti]: our key components (deformable triplane decoder, CVAE formulation, and proposed stage 1 model) each contribute to significant improvements in semantic estimation accuracy.
@@ IV-C Visualizations
Qualitative Results: Figure 4 qualitatively compares our approach with VoxFormer [li2023voxformer] and MonoScene [cao2022monoscene]. Blue rectangles highlight areas present in the Ground Truth but missed by VoxFormer and MonoScene. Red rectangles indicate regions where VoxFormer and MonoScene made incorrect semantic estimations. VoxFormer shows considerable noise in its predictions, especially outside the camera’s FOV, and struggles with occluded areas (e.g., the driving road in the third row). MonoScene has difficulty estimating semantics in occluded areas and completing the map. In contrast, our ET-Former effectively handles occluded and out-of-FOV areas.
@@ V Conclusion, Limitations, and Future Work
ET-Former offers an efficient and robust solution for 3D semantic scene completion using a single monocular camera. It successfully achieves state-of-the-art performance with much less memory usage, highlighting its potential for real-world applications in autonomous navigation systems. Our future work could focus on designing a more memory-efficient yet effective encoder. Additionally, we aim to design a navigation strategy that makes safe but also permissible decisions leveraging both semantic and uncertainty maps generated from the proposed algorithm.